OpenStudy (anonymous):

x-5=square root of 31-9x

OpenStudy (lalaly):

(x-5)^2 =31-9x x^2 -10x +25=31-9x x^2 -x-6=0 (x-3)(x+2)=0 x=3 or x=-2

OpenStudy (anonymous):

tnx so much!